Determine the month and day when the Sun is at minimum altitude for an observer in Washington, D.C. (39°N). What is the altitude of the Sun at this time?
Answer #1

sun will have minimum altitude in winter solstice. Probably, on 22nd December

On this day, sun is at an angle 23.5

so,

zenith = 39 - 23. 5 = 15.5o

Altitude which sun reaches on this day is

90 - 39 - 23.5 = 27.5o